# Focus On This, Not That! Steering LLMs With Adaptive Feature Specification

### v1

@@ 4.1 Validation of FIT on the SS dataset
Figure 3: SS focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on the SS dataset. Here, $C$ refers to the causal feature (sentiment) and $S$ the spurious feature (presence of one of the keywords of “Bayesian” and “Pineapple”).
@@
Figure 4: SMNLI focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on the SMNLI dataset. Here, $C$ refers to the causal feature (logical relationship between premise and hypothesis) and $S$ the spurious feature (genre of the underlying text)
@@
Figure 5 shows the focus accuracy of models on the distribution-shifted test sets across different values of $\rho_{\text{spurious}}$. When focusing on the causal feature or ignoring the spurious feature, the model maintains strong performance in terms of focus accuracy, even on unseen genre values (over $80\%$ focus accuracy for FIT models on the second row of Figure 5). Note that, while we observe low focus accuracy when focusing on spurious features, this is expected, as the spurious labels associated with these new genres were not encountered during training. Thus when focusing on these features the model does not know what label to predict. This result highlights that the focus-tuned models have indeed learned spurious associations during training and correctly reproduces them when instructed to focus on these spurious features, even for new spurious feature values. When instructed to focus on the causal feature (or even just to ignore the spurious feature), the model still shows strong generalisation in the presence of distribution shift.
@@
Figure 5: SMNLI focus accuracies ($\uparrow$) under distribution shift . Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on SMNLI, evaluated on distribution-shifted test sets with different feature values (genres) from training. $\text{shift}(\rho_{\text{spurious}})$ refers to test sets where feature values (genres) differ from training with a co-occurrence rate of $\rho_{\text{spurious}}$. Here, “flipped” indicates a change in spurious label associations between training and testing, with high $\rho_{\text{spurious}}$ in the test set.
@@ 6 Conclusions
In this work, we introduce Focus Instruction Tuning (FIT), a method designed to steer the behaviour of LLMs by focusing on or ignoring specific features when formulating responses. Across a range of tasks and settings, we demonstrate that FIT can be used to steer LLM behaviours at inference time, even in the context of distribution shifts over feature values or when generalizing to unseen features at inference time. Additionally, we show that our method can mitigate biases by identifying and factoring out known stereotypes that might otherwise influence responses. Thus, FIT represents a step toward enabling more robust, fair, and controllable LLMs.
We recommend that future work explore the effectiveness of FIT across a broader variety of tasks, including open-ended, free-form natural language generation tasks such as summarization or translation. Another promising direction is investigating whether FIT can generalise not only across features but also across different categories of tasks (cf. FLAN; Longpre et al. 2023).
Ethical Considerations and Limitations
The ability to steer model behaviours by dynamically focusing on or ignoring arbitrary features such as demographic categories, as enabled by FIT, carries the potential to reduce algorithmic discrimination and mitigate related harms. Indeed, practitioners can measure the discrepancy in behaviour when a model focuses or ignores some features in order to identify and correct discriminatory behaviour. Furthermore, the ability to attribute model responses to specific input features (as enabled by, e.g., observing changes in model behaviour resulting from different focus specifications) is also an important goal in explainable AI, indicating a potential role for FIT in facilitating more direct and productive interactions between humans and AI systems in collaborative decision processes, where insights about the features that caused a model to make a particular prediction may be helpful for assessing whether or not this prediction is an ethical, legal, or responsible basis for a real-world decision. Finally, the ability to focus on user-specified features carries implications for robustness, as focus-tuned models can be steered to focus on stable (causal) features whose relationships with labels are expected to hold across multiple test domains while ignoring (spurious) features whose associations with labels may not generalise beyond the training domain.
On the other hand, one possible negative impact of FIT is that it could potentially be used by bad actors to intentionally bias models to, e.g., make discriminatory predictions based on social bias features. However, this opportunity would already be available by directly fine-tuning the model on a biased dataset. Another possible concern raised by our findings reported in Section A.9 is that FIT may not be as effective when focusing on or ignoring features that heavily overlap with each other. While this may be consideration in certain narrow contexts, it does not detract from FIT’s general promise across arbitrary natural-language tasks (see Section A.9 for further discussion).
@@ A.1 FT Training and Optimisation Settings
Choice of $\rho_{\text{spurious}}$ during training. Consider a classification problem with $N$ classes so that $|\mathcal{Y}|=6$. During FIT training we want the model to learn to change it’s behaviour depending on which features are specified to be focused on or ignored.
To achieve this, consider when the focus instruction requires that the model focus on a spurious feature during training, which means that $I_{\text{focus}}\in\{\text{focus}(S),\;\text{focus}(S)\wedge\text{ignore}(F_{j})\mid F_{j}\in\mathcal{F}\setminus\{S\}\}$, we choose $\rho_{\text{spurious}}=1/N.$ The definition of $\rho_{\text{spurious}}$ given in Equation 4 implies that in this setting $\rho_{\text{spurious}}=\mathbb{P}(Y=y_{s}|X,I,I_{\text{focus}},s\in X)$ for a feature value $s$ of spurious feature $S$ in input $X$. Therefore, setting $\rho_{\text{spurious}}=1/N$ for the training dataset induces a uniform distribution over the set the set of class labels conditioned on an input $X$ and focus instruction $I_{\text{focus}}$ during training.
The entropy of this discrete distribution is then given by:
$\mathcal{H}(\mathbb{P}(Y=y_{s}|X,I,$ $I_{\text{focus}},s\in X))=$ (6)
$=\mathbb{E}_{y\sim\mathbb{P}(Y=y_{s}|X,I,I_{\text{focus}},s\in X)}\left[-\log\mathbb{P}(Y=y_{s}|X,I,I_{\text{focus}},s\in X)\right]$ (7)
$=-\sum_{y\in\mathcal{Y}}\mathbb{P}(Y=y_{s}|X,I,I_{\text{focus}},s\in X)\log\mathbb{P}(Y=y_{s}|X,I,I_{\text{focus}},s\in X)$ (8)
$=-\sum_{y\in\mathcal{Y}}\log\left(\frac{1}{N}\right)\frac{1}{N}$ (9)
$=-\log\left(\frac{1}{N}\right)$ (10)
$=\log N.$ (11)
It is well known that discrete uniform distributions have maximum entropy (Bishop & Nasrabadi, 2006).
Practically, this means that the causal feature is not predictive of the focus label, which is what we are training the model to predict. A lower-entropy distribution (compared to the uniform distribution) during training would result in a higher co-occurrence of the spurious labels with the underlying causal labels. This could lead the model to rely on the underlying task-causal feature to solve the task, rat her than learning to adapt its behaviour when focusing on a non-causal feature. Therefore, using the maximum-entropy uniform distribution during training better enables the model to learn to adjust its behaviour based on the specified features. This ensures that the model does not fall back on causal features when spurious ones are the focus, thus improving the steerability of the model.
A.2 Evaluation Metrics
@@
A.3 FIT Focus Instructions and Prompt Templates
@@
Attention Prompt Template <INSTRUCTION>
@@
A.4 Baselines
SFT baseline. We implement an SFT baseline that follows the same training procedure as FIT, except during training, we exclude any focus instructions from the input prompts while still training on the focus labels. This provides a fair comparison with FIT, as the models are trained on the same input text and label pairs. The rest of the training setup, including hyperparameters and early stopping, remains identical to the FIT training setup.
Few-shot baseline. This second baseline compares FIT training to few-shot inference using the original pre-trained models without additional fine-tuning on our spurious datasets. Specifically, we use 5 in-context examples across all datasets. For the in-context examples, we concatenate multiple examples one after the other, including the instructional prompt only for the first in-context example and the final test example. Each in-context example contains the same focus instruction as the test example for which they serve as context.
A.5 SMNLI Ablation of Training and Test time Focus Instruction Rephrasing Differences
@@
Figure 7: Focus accuracy ($\uparrow$) for different training and test focus instruction sets. Figure comparing focus accuracies ($\mathcal{A}_{\text{focus}}$) of sampling from the same (left) and different (right) sets of focus instructions at training and test time of models on the SMNLI dataset.
A.6 Spurious Sentiment (SS) Dataset
@@
  • $C$ true underlying sentiment, the causal feature within this task, sampled from the original dataset.
  • $S$ spurious feature, here this is the presence of the keywords Bayesian or Pineapple. We represent this as a binary vector $S\in\{0,1\}^{2}$, where the first and second components of this vector denote the presence of either the keyword Pineapple or Bayesian respectively.
@@
$S$ $=f_{S}(C,U_{S});$ (14)
$X$ $=f_{X}(\tilde{X},S,U_{X},U_{\text{incld.}});$ (15)
@@
$S=\left(\bm{1}_{C=0},\bm{1}_{C=1}\right);$ (19)
@@
$U_{X}\sim\texttt{Ber}(\rho_{\text{spurious}});$ (21)
$X=\begin{cases}U_{X}\texttt{LLM}(\tilde{X},S)+(1-U_{X})\texttt{LLM}(\tilde{X},1-S)&\text{if }U_{\text{incd.}}=1\\ \tilde{X}&\text{ if }U_{\text{incd.}}=0\end{cases};$ (22)
@@
The variable $\rho_{C}$ gives the distribution of sentiment labels in the original binarised SST-5 dataset. Moreover, $p_{\mathcal{D}}(\cdot|C)$ denotes the conditional dataset distribution of the different input texts give $C$ (here we assume that we are just uniformly sampling text with the given sentiment $C$) and $\bm{1}_{(\cdot)}$ denotes the indicator function. In addition, $\rho_{\text{incl.}}$ gives the proportion of spurious features that are included within original dataset examples. This corresponds to proportion of examples within the original dataset that are modified by the above process and therefore contain spurious feature values.
Finally, we prove that $\rho_{\text{spurious}}$ gives the concurrence rate between the label $Y$ and the spurious feature values of $S$. The proof rests on the fact that $U_{X}$, which gives whether a spurious feature value $s$ is injected into $\tilde{X}$ or whether the other value $1-s$ is injected instead, controls the cooccurrence between $Y$ and the spurious feature value $s$. In particular, we note that $\mathbb{P}(Y=y,U_{X}=1,X\neq\tilde{X}|\tilde{X},S=s)$ gives the co-occurrence rate between the label $y$ and spurious feature $s$ in the dataset (assuming that if the feature is only present within the dataset through the inclusion of the feature and not within the original dataset examples).
@@
From the DGP described in Figure 8, we have that
$\mathbb{P}(Y=1,U_{X}=1,X\neq\tilde{X}|\tilde{X},S=(0,1))=\rho_{\text{incdl.}}\cdot\rho_{\text{spurious}};$ (24)
$\mathbb{P}(Y=0,U_{X}=1,X\neq\tilde{X}|\tilde{X},S=(1,0))=\rho_{\text{incdl.}}\cdot\rho_{\text{spurious}}.$ (25)
@@
Using the chain rule of probability, we see that
$\mathbb{P}($ $Y=1,U_{X}=1,X\neq\tilde{X}|\tilde{X},S=(0,1))=$ (26)
$=\mathbb{P}(Y=1|U_{X}=1,\tilde{X},X\neq\tilde{X},S=(0,1))\mathbb{P}(U_{X}=1,X\neq\tilde{X}|\tilde{X},S=(0,1))$ (27)
$=\mathbb{P}(Y=1|S=(0,1))\mathbb{P}(U_{X}=1,X\neq\tilde{X}|\tilde{X},S=(0,1))$ (28)
$=\mathbb{P}(Y=1|S=(0,1))\mathbb{P}(U_{X}=1|\tilde{X},X\neq\tilde{X},S=(0,1))\mathbb{P}(X\neq\tilde{X}|\tilde{X},S=(0,1))$ (29)
$=\mathbb{P}(Y=1|S=(0,1))\mathbb{P}(U_{X}=1)\mathbb{P}(X\neq\tilde{X}|\tilde{X})$ (30)
$=\mathbb{P}(Y=1|C=1)\mathbb{P}(U_{X}=1)\mathbb{P}(U_{\text{incld.}}=1)$ (31)
$=1\cdot\rho_{\text{incdl.}}\cdot\rho_{\text{spurious}}$ (32)
$=\rho_{\text{incdl.}}\cdot\rho_{\text{spurious}},$ (33)
where we have used that $S$ and $C$ share a deterministic relationship and have used the conditional independencies specified within the graphical model depicted in Figure 8, and through the noise terms $U_{(\cdot)}$ in the SCEs introduced above. ∎
With $\rho_{\text{incld.}}=1$, we have that, indeed, the co-occurrence rate between the presence of spurious feature values and the labels $Y$ is given by $\rho_{\text{spurious}}$.
We now connect this to predicitvity which is defined Equation 4 in order to justify the notation that we are using within the SCEs above.
Corollary 1.
From the DGP described in Figure 10, we have that
$\mathbb{P}(Y=1|\tilde{X},S=(0,1),U_{X}=1,X\neq\tilde{X})=\rho_{\text{spurious}};$ (34)
$\mathbb{P}(Y=0|\tilde{X},S=(1,0),U_{X}=1,X\neq\tilde{X})=\rho_{\text{spurious}}.$ (35)
These events correspond to the events given in Equation 4 that define $\rho_{\text{spurious}}(s)$ generally. Therefore, the notation of $\rho_{\text{spurious}}$ in Equation 4 is justified.
@@
Using Proposition 1 and the definition of conditional probability, we get that
$\mathbb{P}(Y=1\mid\tilde{X},S=(0,1),$ $U_{X}=1,X\neq\tilde{X})$ (36)
$=\frac{\mathbb{P}(Y=1,U_{X}=1,X\neq\tilde{X}\mid\tilde{X},S=(0,1))}{\mathbb{P}(U_{X}=1,X\neq\tilde{X}\mid\tilde{X},S=(0,1))}$ (37)
$=\frac{\mathbb{P}(Y=1,U_{X}=1,X\neq\tilde{X}\mid\tilde{X},S=(0,1))}{\mathbb{P}(U_{X}=1\mid\tilde{X},S=(0,1),X\neq\tilde{X})\mathbb{P}(X\neq\tilde{X}|\tilde{X},S=(0,1))}$ (38)
$=\frac{\mathbb{P}(Y=1,U_{X}=1,X\neq\tilde{X}\mid\tilde{X},S=(0,1))}{\mathbb{P}(U_{X}=1)\mathbb{P}(X\neq\tilde{X}|\tilde{X},U_{X}=1)}$ (39)
$=\frac{\rho_{\text{icld.}}\cdot\rho_{\text{spurious}}}{\rho_{\text{icld.}}\cdot 1}$ (40)
$=\rho_{\text{spurious}},$ (41)
where we have used the independencies specified within the DGP in Figure 10 and the noise term $U_{x}$ within the functional description of this DGP. ∎

### v2

@@ 4.1 Validation of FIT on the SS dataset
Figure 3: SS focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on the SS dataset. Here, $C$ refers to the core feature (sentiment) and $S$ the spurious feature (presence of one of the keywords of “Bayesian” and “Pineapple”).
@@
Figure 4: SMNLI focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on the SMNLI dataset. Here, $C$ refers to the core feature (logical relationship between premise and hypothesis) and $S$ the spurious feature (genre of the underlying text)
@@
Figure 5 shows the focus accuracy of models on the distribution-shifted test sets across different values of $\rho_{\text{spurious}}$. When focusing on the core feature or ignoring the spurious feature, the model maintains strong performance in terms of focus accuracy, even on unseen genre values (over $80\%$ focus accuracy for FIT models on the second row of Figure 5). Note that, while we observe low focus accuracy when focusing on spurious features, this is expected, as the spurious labels associated with these new genres were not encountered during training. Thus when focusing on these features the model does not know what label to predict. This result highlights that the focus-tuned models have indeed learned spurious associations during training and correctly reproduces them when instructed to focus on these spurious features, even for new spurious feature values. When instructed to focus on the core feature (or even just to ignore the spurious feature), the model still shows strong generalisation in the presence of distribution shift.
@@
Figure 5: SMNLI focus accuracies ($\uparrow$) under distribution shift . Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on SMNLI, evaluated on distribution-shifted test sets with different feature values (genres) from training.
@@ 6 Conclusions
In this work, we introduce Focus Instruction Tuning (FIT), a method designed to steer the behaviour of LLMs by focusing on or ignoring specific features when formulating responses. Across a range of tasks and settings, we demonstrate that FIT provides dynamic and precise control over LLM behaviour at inference time, enabling users to adapt model responses even in the context of distribution shifts over feature values or when generalising to unseen features. Furthermore, our approach can address challenges such as mitigating the influence of known stereotypes that might otherwise impact responses, showcasing one of its many applications. Thus, FIT represents a step toward enabling more robust, steerable, fair, and controllable LLMs.
We recommend future work explore FIT’s effectiveness in a broader range of tasks, including open-ended generation tasks like summarization or translation, which were not covered here due to the additional costs and challenges associated with curating high-quality datasets for open-ended NLG tasks. Another direction is investigating FIT’s ability to generalise across task categories.
Ethical Considerations
The ability to dynamically steer model behaviour by focusing on or ignoring features, as enabled by FIT, holds significant potential for reducing algorithmic discrimination and mitigating harms. Practitioners can leverage FIT to identify and correct biases by measuring discrepancies in behaviour when a model focuses on or ignores specific features. Additionally, FIT enhances explainability by attributing model predictions to input features, enabling more transparent and productive human-AI collaboration. This supports ethical and responsible decision-making by assessing whether predictions are justified. FIT also enhances robustness by prioritising stable core features expected to generalise across domains while ignoring spurious, domain-specific biases, making it a valuable tool for fairness, explainability, and robustness. However, risks include potential misuse by bad actors to bias models, though this is not unique to FIT and could already be achieved through biased fine-tuning.
@@ A.1 FT Training and Optimisation Settings
Choice of $\rho_{\text{spurious}}$ during training. Ensuring the independence $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training is critical for inducing steerability through feature specification in the FIT framework.
  • Independence $Y\perp\!\!\!\perp S$: prevents the model from leveraging the spurious feature $S$ to predict the ground-truth label $Y$. With no predictive signal flowing from $S$ to $Y$, the model must rely exclusively on the core feature $C$ for accurate label predictions. This design choice safeguards the model from overfitting to spurious correlations, thereby maintaining robust performance under distribution shifts. Moreover, removing any inherent relationship between $S$ and $Y$ ensures that for focus instruction intending for the model to utilise the core feature $C$ only during inference, the model cannot exploit a potential shortcut using $S$; it must utilise the core feature alone for prediction in this scenario enabling prediction only through the specified feature indicated through the focus instruction passed to the model.
  • Independence $Y_{S}\perp\!\!\!\perp C$: serves a complementary role of preventing the model from exploiting the core feature $C$ when predicting spurious labels $Y_{S}$. By ensuring $C$ carries no information about $Y_{S}$, the model cannot use the true task feature $C$ as a shortcut for spurious-label predictions; it must again learn to only use the specified feature within the passed focus instructions alone for making predictions.
When either independence condition is violated (e.g., $Y\mathrel{\ooalign{$\perp\!\!\!\perp$\cr\hss$/$\hss}}S$ or $Y_{S}\perp\!\!\!\perp C$), the model can exploit correlations that undermine the FIT objective. For instance, if $Y\mathrel{\ooalign{$\perp\!\!\!\perp$\cr\hss$/$\hss}}S$, instructing the model to $\text{ignore}(S)$ would not necessarily prevent it from using $S$ to predict $Y$. This failure to follow instructions defeats the purpose of feature-based steerability. Thus, both $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ are necessary for building a training environment in which the model learns to precisely follow focus instructions and adjust its behaviour in accordance with specified features.
To achieve this independence in our synthetic SS and SMNLI datasets, we set $\rho_{\text{spurious}}=1/N$, where $N$ is the number of class labels. Additionally, we enforce a balanced label distribution in the training set to eliminate any indirect biases that could correlate $S$ with $Y$. As shown in Section A.9 and Section A.11, these conditions are sufficient to guarantee $Y\perp\!\!\!\perp S$ in the training data, enabling the model to effectively learn steerable behaviour from focus instructions.
A.2 Limitations and Future Work
While Focus Instruction Tuning (FIT) relies on some prior knowledge of spurious features during training, this requirement aligns with existing practices in industry and research (Deldjoo, 2023; Microsoft, 2020), as well as current regulatory and ethical expectations (Guldimann et al., 2024; Zeng et al., 2024). Moreover, pre-trained models can generate focus labels automatically when explicit spurious feature labels are unavailable. For example, our method is compatible with existing and future approaches for automatically identifying and labeling spurious features (Wang & Culotta, 2020; Zhou et al., 2023; Wang et al., 2022; Zheng et al., 2024). FIT can act as a complementary mechanism, providing steerability based on identified features, whether they are manually labeled or automatically detected.
Our experiments primarily focus on classification and multiple-choice QA datasets due to the cost and challenges associated with curating high-quality datasets for open-ended NLG tasks. However, this reflects a pragmatic prioritisation of introducing a novel methodology over exhaustive data collection, rather than a limitation of FIT itself. Extending FIT to open-ended tasks, such as summarization or translation, remains an exciting direction for future research, as does exploring its ability to generalise across diverse task categories using setups similar to FLAN (Longpre et al., 2023).
Additionally, our evaluation on the HANS dataset revealed challenges when addressing overlapping or less-distinctive features. While FIT demonstrated strong performance in generalising and steering models based on identified features, overlapping heuristics can introduce ambiguity, highlighting the need for further refinements in handling such cases. Despite these limitations, FIT represents a promising foundation for enabling more robust, fair, and controllable LLMs across a range of tasks.
A.3 Evaluation Metrics
@@
A.4 FIT Focus Instructions and Prompt Templates
@@
BBQ focus instruction prompt template $\mathcal{I}_{\text{focus}}$ <INSTRUCTION>
@@
A.5 Baselines
We include results for the following two baselines to fruther supplement the results presented in the main section of the paper.
SFT($y_{\text{focus}}$) baseline. We implement an SFT baseline that follows the same training procedure as FIT, except during training, we exclude any focus instructions from the input prompts while still training on the focus labels. This provides a fair comparison with FIT, as the models are trained on the same input text and label pairs. The rest of the training setup, including hyperparameters and early stopping, remains identical to the FIT training setup.The model is tested on the full set of focus instructions prompts detailed in Equation 1.
Few-shot baseline. This second baseline compares FIT training to few-shot inference using the original pre-trained models without additional fine-tuning on our spurious datasets. Specifically, we use 5 in-context examples across all datasets. For the in-context examples, we concatenate multiple examples one after the other, including the instructional prompt only for the first in-context example and the final test example. Each in-context example contains the same focus instruction as the test example for which they serve as context. The model is tested on the full set of focus instructions prompts detailed in Equation 1.
For completeness, we report two additional baselines: vanilla SFT and zero-shot baselines.
SFT($y$) baseline. We implement a vanilla SFT baseline that simply trains a model using SFT on inputs and their ground truth labels (as opposed to focus labels in the SFT($y_{\text{focus}}$) baseline). During training, only standard IT prompts are used, with no additional focus instructions included. The rest of the training setup, including hyperparameters and early stopping, remains identical to the FIT training setup. The model is tested on the full set of focus instructions prompts detailed in Equation 1.
Zero-shot baseline. Finally, we include a zero-shot inference baseline using the original pre-trained models without additional fine-tuning on our spurious datasets. No in-context examples are used at inference time, and the model is not trained at all beyond it’s pre-training. The model is tested on the full set of focus instructions prompts detailed in Equation 1.
A.6 Additional Baselines Results
In this section, we include the two additional baselines -SFT($y$) and zero-shot - on the SMNLI dataset to further supplement the results in Section 4.2.
Figure 7: Full Baseline vs FIT Focus accuracy ($\uparrow$) on SS. Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the SS dataset.
Figure 8: Full Baseline vs FIT Focus accuracy ($\uparrow$) on SMNLI. Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the SMNLI dataset.
Figure 9: Full Baseline vs FIT Focus accuracy ($\uparrow$) on BBQ. Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the BBQ dataset.
A.7 SMNLI Ablation of Training and Test time Focus Instruction Rephrasing Differences
@@
Figure 10: Focus accuracy ($\uparrow$) for different training and test focus instruction sets. Figure comparing focus accuracies ($\mathcal{A}_{\text{focus}}$) of sampling from the same (top) and different (bottom) sets of focus instructions at training and test time of models on the SMNLI dataset.
@@
A.8 Comparison of FIT against a specific debiasing technique
FIT is a general framework designed to enable users to steer a model’s behavior based on specified features. This approach provides enhanced control over model outputs during inference, adding a critical layer of explainability and controllability to model predictions.
While understanding and mitigating biases or spurious correlations is a valuable and natural application of FIT, it is not the sole objective. The broader goal of steerability includes addressing challenges in managing and aligning model behaviour across diverse contexts. For instance, maintaining controllability is crucial in addressing safety alignment fragility, which can emerge after fine-tuning (Bhattacharjee et al., 2024). In such cases, the ability to adapt model responses to align with user specifications ensures safe and reliable deployment.
Experiment. To explore FIT’s broader applicability, we compare its performance as a debiasing method against a well-known debiasing technique: the Product of Experts (PoE) method (Mahabadi et al., 2020). PoE involves training a bias model $f_{B}$, which is trained exclusively on bias features. This bias model mediates the training of the final model $f$ by combining their predictions through an elementwise product: $\sigma(f(x))\odot\sigma(f_{B}(x_{B}))$, where $x\in\mathcal{D}$ and $x_{B}$ represents the biased feature of $x$.
We adapted this approach to our setting by training a bias model on the stereotypical labels within the BBQ dataset. These labels correspond to group-stereotypical associations. For autoregressive models, we further modified the PoE method by extracting and normalizing the logits of the first newly generated token position over the set of single tokens representing the answer options.
Results. The results of the debiasing experiment comparing FIT to the PoE method is shown in Figure 11. FIT performs equally as well as the PoE method as shown by the comparing the default prompt accuracy () for the PoE models against the $\text{focus}(C)$ results for the FIT models; both metrics correspond to causal accuracy for these prompt types. Indicating that FIT performs just as well as a dedicated debiasing technique.
However, the PoE method requires training two separate models and does not provide steerability at test time as shown by the low focus accuracy on $\text{focus}(S)$. Indeed the model defaults to the ground truth label across all prompt types and does not change behaviour despite different different focus specifications. This highlights the flexibility of FIT, which not only debiases effectively but also enables additional controllability during inference.
Figure 11: Focus Accuracy ($\uparrow$) of FIT against PoE Debiasing Technique. Figure showing the focus accuracies ($\mathcal{A}_{\text{focus}}$) of FIT (bottom row) and the dedicated debaiasing technique, PoE (top row), on the BBQ dataset.
A.9 Spurious Sentiment (SS) Dataset
@@
  • $C$ true underlying sentiment, the core feature within this task, sampled from the original dataset.
  • $\tilde{S}$ proposed spurious feature sample, here this is the presence of the keywords Bayesian or Pineapple. We represent this as a binary vector $S\in\{0,1\}^{2}$, where the first and second components of this vector denote the presence of either the keyword Pineapple or Bayesian respectively. We restrict to consider only one keyword appearing in a given text at a time so that $\text{Val}(S)=\{(1,0),(0,1)\}$.
  • $S$ final spurious feature that is naturally inserted using a LLM into the final SS dataset example $X$. $S$ is a randomly flipped version of the proposed spurious feature $S$.
@@
$\tilde{S}$ $=f_{\tilde{S}}(C,U_{\tilde{S}});$ (8)
$S$ $=f_{S}(\tilde{S},U_{S});$ (9)
$X$ $=f_{X}(\tilde{X},S,U_{X});$ (10)
@@
$\tilde{S}=\left(\bm{1}_{C=0},\bm{1}_{C=1}\right);$ (14)
$U_{S}\sim\texttt{Ber}(\rho_{\text{spurious}});$ (15)
$S=U_{S}\tilde{S}+(1-U_{S})(1-\tilde{S});$ (16)
@@
$X=\texttt{LLM}(\tilde{X},S);$ (18)
@@
The variable $\rho_{C}$ gives the distribution of sentiment labels in the original binarised SST-5 dataset. Moreover, $p_{\mathcal{D}}(\tilde{x}|C)$ denotes the conditional dataset distribution of the different input texts give $C$ (here we assume that we are just uniformly sampling text with the given sentiment $C$) and $\bm{1}_{(\cdot)}$ denotes the indicator function.
Finally, we prove that $\rho_{\text{spurious}}$ gives the coocurrence rate/predictivity between the label $Y$ and the spurious feature $S$, and is well-defined notation in the sense that it corresponds to Definition 1 so that $\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s)$, where $y_{s}$ is the label that spurious feature value $s$ is spuriously correlated with.
@@
From the SCEs described above, assuming that we have a balanced label distribution, that is $\rho_{C}=1/2$, we have that
$\mathbb{P}(Y=y_{s}\mid S=s)=\rho_{\text{spurious}}\;.$ (20)
for all spurious feature values $s$.
@@
First note that we have by Equation 12 and the assumption that $\rho_{C}=1/2$, that $\mathbb{P}(Y=1)=\mathbb{P}(Y=0)=1/2$. Moreover, using the partition theorem, we have that
$\mathbb{P}(S=s)$ $=\mathbb{P}(S=s\mid\tilde{S}=s)\mathbb{P}(\tilde{S}=s)+\mathbb{P}(S=s\mid\tilde{S}\neq s)\mathbb{P}(\tilde{S}\neq s)$ (21)
$=\rho_{\text{spurious}}\cdot\frac{1}{2}+(1-\rho_{\text{spurious}})\cdot\frac{1}{2}$ (22)
$=\frac{1}{2}.$ (23)
Finally note that $\mathbb{P}(S=s\mid Y=y_{s})=\rho_{\text{spurious}}$ as $Y=y_{s}$ forces $C=y_{s}$, and therefore that $\tilde{S}=s$ by construction.
Putting this all together and utilising Bayes’ rule gives
$\mathbb{P}(Y=y_{s}\mid S=s)$ $=\frac{\mathbb{P}(S=s\mid Y=y_{s})\mathbb{P}(Y=y_{s})}{\mathbb{P}(S=s)}$ (24)
$=\frac{\mathbb{P}(S=s\mid Y=y_{s})\mathbb{P}(Y=y_{s})}{\mathbb{P}(S=s\mid\tilde{S}=s)\mathbb{P}(\tilde{S}=s)+\mathbb{P}(S=s\mid\tilde{S}\neq s)\mathbb{P}(\tilde{S}\neq s)}$ (25)
$=\frac{\rho_{\text{spurious}}\cdot\frac{1}{2}}{\frac{1}{2}}$ (26)
$=\rho_{\text{spurious}}.$ (27)
which gives the result. ∎
Within our experiments in Section 4.1, we always force the label distribution to be balanced, that is $\rho_{C}=1/2$, and assume that within each dataset split, $\rho_{\text{spurious}}$ is the same rate for all spurious feature values.
Independence conditions during training for FIT. As specified in Section A.1, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features. Here, $Y_{S}$ is the spurious label spuriously correlated to spurious feature value $S$. The results below give sufficient conditions for these independence conditions to be specified with respect to the DGP described in Figure 12.
Proposition 2.
Assuming the SCEs given above and the corresponding DGP described in Figure 12, if we have a uniform label distribution $p(Y=y)$, that is $\rho_{C}=1/2$, and have that $\rho_{\text{spurious}}=1/2$ in the SS training set, then we have that $Y\perp\!\!\!\perp S$.
@@
[…]
Proposition 3.
Assuming the SCEs given above and the corresponding DGP described in Figure 12, if we have a uniform label distribution $p(Y=y)$, that is $\rho_{C}=1/2$, and have that $\rho_{\text{spurious}}=1/2$ in the SS training set, then we have that $Y_{S}\perp\!\!\!\perp C$.
Proof.
[…]
$\mathbb{P}(S=s\mid C=c)$ $=\sum_{\tilde{s}}\mathbb{P}(S=s\mid C=c,\tilde{S}=\tilde{s})\mathbb{P}(\tilde{S}=\tilde{s}\mid C=c)$ (28)
$=\sum_{\tilde{s}}\mathbb{P}(S=s\mid\tilde{S}=\tilde{s})\mathbb{P}(\tilde{S}=\tilde{s}\mid C=c)$ (29)
$=\underbrace{\mathbb{P}(S=s\mid\tilde{S}=s)}_{=\rho_{\text{spurious}}}\underbrace{\mathbb{P}(\tilde{S}=s\mid C=c)}_{=\frac{1}{2}}$ (30)
$+\underbrace{\mathbb{P}(S=s\mid\tilde{S}\neq s)}_{=1-\rho_{\text{spurious}}}\underbrace{\mathbb{P}(\tilde{S}\neq s\mid C=c)}_{\frac{1}{2}}$ (31)
$=\frac{1}{2}.$ (32)
[…]